# Kernel Mean Embedding Topology: Weak and Strong Forms for Stochastic Kernels and Implications for Model Learning

### v1

@@ 3.2 Two weak topologies revisited
3.2.1 $w^{*}$-topology on Stochastic Kernels
In this section, we introduce the $w^{*}$-topology on the set of stochastic kernels, a concept originally developed in [11, Section 2.4] for randomized Markov policies to prove the existence of optimal policies for continuous-time stochastic control problems with average cost optimality criterion. Later, in [51], $w^{*}$-topology is used to establish the existence of optimal policies in team decision problems. This topology has proven to be a valuable tool in analyzing such problems. We will begin by recalling its formal definition and then proceed to discuss some of its fundamental properties.
@@ 3.2.1 $w^{*}$-topology on Stochastic Kernels
which turns $(C_{0}({\mathsf{U}}),\|\cdot\|)$ into a separable Banach space. Let $\|\cdot\|_{TV}$ denote the total variation norm on ${\mathcal{M}}({\mathsf{U}})$. The following result is a famous Riesz-Markov-Kakutani representation theorem, which states that the topological dual of $C_{0}({\mathsf{U}})$ is ${\mathcal{M}}({\mathsf{U}})$.
@@
[25, Theorem 7.17] For any $\nu\in{\mathcal{M}}({\mathsf{U}})$ and $g\in C_{0}({\mathsf{U}})$, let $I_{\nu}(g)\triangleq\langle\nu,g\rangle$, where
@@
Furthermore, the norm on ${\mathcal{M}}({\mathsf{U}})$ induced by this duality is total variation norm.
@@
$\|\gamma\|_{q}$ =$=\begin{cases}\left(\int_{{\mathsf{Y}}}\|\gamma(y)\|^{q}_{TV}\,\mu(dy)\right)^{1/q},&\,\,\text{if}\,\,q<\infty\\ {\mathrm{ess}}\sup_{y\in{\mathsf{Y}}}\|\gamma(y)\|_{TV},&\,\,\text{if}\,\,q=\infty.\end{cases}$ (6)
Since $C_{0}({\mathsf{U}})$ is separable, $L_{p}\bigl{(}\mu,C_{0}({\mathsf{U}})\bigr{)}$ is also separable. By Theorem 2.1, the topological dual of $L_{p}\bigl{(}\mu,C_{0}({\mathsf{U}})\bigr{)}$ is $\mathcal{L}_{q}\bigl{(}\mu,{\mathcal{M}}({\mathsf{U}})\bigr{)}$; that is
@@
[29, Definition C.1] A mapping $\gamma:{\mathsf{Y}}\rightarrow{\mathcal{M}}({\mathsf{U}})$ is called stochastic kernel from ${\mathsf{Y}}$ to ${\mathsf{U}}$ if, for all $D\in{\mathcal{U}}$, the mapping ${\mathsf{Y}}\ni y\mapsto\gamma(y)(D)\in\mathbb{R}$ is ${\mathcal{Y}}/{\mathcal{B}}(\mathbb{R})$-measurable and $\gamma(y)\in{\mathcal{P}}({\mathsf{U}})$ for all $y\in{\mathsf{Y}}$. Let $\Gamma$ denote the set of all stochastic kernels from ${\mathsf{Y}}$ to ${\mathsf{U}}$.
@@
Next, consider any closed set $F\subset{\mathsf{U}}$. The indicator function $1_{F}$ can be approximated pointwise by the continuous and bounded functions $h_{n}(u)=\max(1-nd_{{\mathsf{U}}}(u,F),0)$, where $d_{{\mathsf{U}}}$ denotes the metric on ${\mathsf{U}}$ and $d_{{\mathsf{U}}}(u,F)=\inf_{y\in F}d_{{\mathsf{U}}}(u,y)$. Consequently, this implies that the mapping $x\mapsto\gamma(x)(F)\in\mathbb{R}$ is also $\mathcal{X}/\mathcal{B}(\mathbb{R})$-measurable for any closed set $F$ in ${\mathsf{U}}$. Thus, we can conclude that $\gamma$ is a stochastic kernel by [9, Proposition 7.25].
@@
Let ${\mathcal{P}}_{{}_{\leq 1}}({\mathsf{U}})$ denote the set of sub-probability measures in ${\mathcal{M}}({\mathsf{U}})$. Then, since $\mathcal{L}_{q}\bigl{(}\mu,{\mathcal{P}}_{{}_{\leq 1}}({\mathsf{U}})\bigr{)}$ is proved to be closed in $\mathcal{L}_{q}\bigl{(}\mu,{\mathcal{M}}({\mathsf{U}})\bigr{)}$ with respect to $w^{*}$-topology (see [51]) and is a bounded subset, it is (sequentially) compact and metrizable with respect to $w^{*}$-topology by Banach-Alaoglu Theorem. Note that $\mathcal{L}_{q}\bigl{(}\mu,{\mathcal{P}}({\mathsf{U}})\bigr{)}$ is a subset of $\mathcal{L}_{q}\bigl{(}\mu,{\mathcal{P}}_{{}_{\leq 1}}({\mathsf{U}})\bigr{)}$, and so, we can endow $\mathcal{L}_{q}\bigl{(}\mu,{\mathcal{P}}({\mathsf{U}})\bigr{)}$ with relative $w^{*}$-topology. Unfortunately, $\mathcal{L}_{q}\bigl{(}\mu,{\mathcal{P}}({\mathsf{U}})\bigr{)}$ is not closed in this topology as shown below [^5] In [51], we established this result by constructing a counterexample specifically for the case when ${\mathsf{U}}=\mathbb{R}$, which is relatively easy. In Lemma 3.8 above, we extend this by providing a counterexample applicable to any locally compact space ${\mathsf{U}}$. This counterexample draws inspiration from the elegant proof presented in [55, Lemma 12]..
@@
First note that if ${\mathsf{U}}$ is compact, then $C_{0}({\mathsf{U}})=C_{b}({\mathsf{U}})$. Moreover, if $\gamma_{\lambda}\rightharpoonup^{*}\gamma$ with respect to $w^{*}$-topology, then $\mu\otimes\gamma_{\lambda}({\mathsf{Y}}\times\cdot)\rightarrow\mu\otimes\gamma({\mathsf{Y}}\times\cdot)$ in vague topology, and so, in weak convergence topology as $C_{0}({\mathsf{U}})=C_{b}({\mathsf{U}})$. Since $\mu\otimes\gamma_{\lambda}({\mathsf{Y}}\times\cdot)\in{\mathcal{P}}({\mathsf{U}})$ for all $\lambda$ and ${\mathcal{P}}({\mathsf{U}})$ is a closed subset of ${\mathcal{M}}({\mathsf{U}})$ under weak convergence topology, we have $\mu\otimes\gamma({\mathsf{Y}}\times\cdot)\in{\mathcal{P}}({\mathsf{U}})$. Hence, $\gamma\in\mathcal{L}_{q}\bigl{(}\mu,{\mathcal{P}}({\mathsf{U}})\bigr{)}$.
@@
Since $\varepsilon$ is arbitrary, we have $\int_{{\mathsf{U}}}g(u)\,\nu_{n}(du)\rightarrow 0$ as $n\rightarrow\infty$ for any $g\in C_{0}({\mathsf{U}})$. Hence, $\nu_{n}$ converges to 00, where $0(\,\cdot\,)$ denotes the degenerate measure on ${\mathsf{U}}$; that is, $00(D)=$ 0 for all $D\in{\mathcal{U}}$.
@@
[…]
@@ 3.2.2 The Young Narrow Topology on Stochastic Kernels
Conversely, let $h\in\mathop{\rm Car}_{0}({\mathsf{X}}\times{\mathsf{U}})$ and define $f_{h}:{\mathsf{X}}\rightarrow C_{0}({\mathsf{U}})$ as $f_{h}(x)=h(x,\,\cdot\,)$. Then, for any $\gamma\in{\mathcal{M}}({\mathsf{U}})$, the function ${\mathsf{X}}\ni x\mapsto\int_{{\mathsf{U}}}f_{h}(x)(u)\,\gamma(du)\in\mathbb{R}$ is measurable. Hence, $f_{h}$ is weakly measurable. But since $C_{0}({\mathsf{U}})$ is separable, $f_{h}$ is also strongly measurable by [46, Theorem 4.2.4-(c)]. ∎
@@
Let us now define Young narrow topology on $\mathcal{L}_{q}\bigl{(}\mu,{\mathcal{M}}({\mathsf{U}})\bigr{)}$ [46, Definition 4.7.11] [^6] Normally, Young narrow topology is defined only on the set of stochastic kernels $\mathcal{L}_{q}\bigl{(}\mu,{\mathcal{P}}({\mathsf{U}})\bigr{)}$. But, here, we extend this definition to the set of all weakly-measurable functions from ${\mathsf{Y}}$ to ${\mathcal{M}}({\mathsf{U}})$. Moreover, in the original formulation, the topology is induced by the set of $C_{b}$-Caratheodory functions that are in $L_{1}(\mu,C_{b}({\mathsf{U}}))$. Here, we extend this definition to $L_{p}(\mu,C_{b}({\mathsf{U}}))$ for any $1\leq p<\infty$.. Young narrow topology on $\mathcal{L}_{q}\bigl{(}\mu,{\mathcal{M}}({\mathsf{U}})\bigr{)}$ is the smallest topology for which the mapping
@@
In Young narrow topology, one can prove that $\mathcal{L}_{q}\bigl{(}\mu,{\mathcal{P}}({\mathsf{U}})\bigr{)}$ is a closed (as opposed to $w^{*}$-topology and weak kernel mean embedding topology) and metrizable subset of $\mathcal{L}_{q}\bigl{(}\mu,{\mathcal{M}}({\mathsf{U}})\bigr{)}$ [46, Proposition 4.7.14]. However, $\mathcal{L}_{q}\bigl{(}\mu,{\mathcal{P}}({\mathsf{U}})\bigr{)}$ is not relatively (sequentially) compact in this topology unless ${\mathsf{U}}$ is compact because of the lack of the duality result (i.e., The Banach-Alaoglu theorem is not applicable in this context).
@@ 3.3 Comparison of Weak Topologies
Interestingly, to achieve closure in the $w^{*}$-topology and the weak kernel mean embedding topology, a tightness condition is required. This condition effectively prevents probability measures from losing mass at infinity, a necessity highlighted by counterexample in Lemma 3.8. Similarly, for the Young narrow topology to attain relative (sequential) compactness, the same tightness condition must be satisfied, as implied by Prokhorov’s theorem. Consequently, in terms of ensuring the existence of optimal solutions, all three topologies provide comparable benefits.
@@
[…]
@@
Hence, $\mu\otimes\gamma_{\lambda}(E\times\cdot)\rightarrow\mu\otimes\gamma(E\times\cdot)$ in weak MMD topology. Since $\mu\otimes\gamma_{\lambda}(E\times{\mathsf{U}})=\mu\otimes\gamma(E\times{\mathsf{U}})=\mu(E)$, by [55, Lemma 5 and Remark 6], $\mu\otimes\gamma_{\lambda}(E\times\cdot)\rightarrow\mu\otimes\gamma(E\times\cdot)$ in MMD topology. This completes the proof of the forward part.
[…]
@@
Note that for a fixed $E\in\mathcal{Y}$, the total masses of the sub-probability measures $\{\mu\otimes\gamma_{\lambda}(E\times\cdot)\}$ and $\mu\otimes\gamma(E\times\cdot)$ are both equal to $\mu(E)$ (assuming, without loss of generality, that $\mu(E)>0$). As used several times so far, by [55, Lemma 5], the MMD topology, vague topology, and weak convergence topology are equivalent for the set of probability measures. Since the aforementioned sub-probability measures have the same mass as $\mu(E)$, we can normalize them using the constant $\mu(E)$ to convert them into probability measures and apply the equivalence result from [55, Lemma 5]. This completes the proof. ∎
@@
In other words, tightness prevents distributions in ${\mathcal{G}}$ from losing mass at infinity.
@@
where $\mu\otimes\gamma(dy,du)\triangleq\mu(dy)\,\gamma(y)(du)$. One can prove that ${\mathcal{C}}_{{\mathcal{G}}}$ is tight in the sense of the Definition 4.7.20 in [46]. Hence, by [46, Theorem 4.7.22], $\mathcal{L}_{q}(\mu,{\mathcal{G}})$ is relatively (sequentially) compact with respect to the Young narrow topology. Conversely, suppose that $\mathcal{L}_{q}(\mu,{\mathcal{G}})$ is relatively (sequentially) compact with respect to the Young narrow topology and ${\mathsf{U}}$ is complete. This implies that ${\mathcal{G}}$ is also relatively (sequentially) compact with respect to the weak topology. Indeed, let $\{\nu_{n}\}$ be a sequence in ${\mathcal{G}}$ and for each $n$, define the constant function $\gamma_{n}(y)=\nu_{n}$ for all $y\in{\mathsf{Y}}$. Since $\mathcal{L}_{q}(\mu,{\mathcal{G}})$ is relatively (sequentially) compact with respect to the Young narrow topology, there exists a subsequence $\{\gamma_{n_{k}}\}$ of $\{\gamma_{n}\}$ that converges with respect to the Young narrow topology. This implies that subsequence $\{\nu_{n_{k}}\}$ of $\{\nu_{n}\}$ also converges with respect to the weak topology. Hence, ${\mathcal{G}}$ is relatively (sequentially) compact with respect to the weak topology. By Prohorov’s theorem [30, Theorem 1.4.12], ${\mathcal{G}}$ is tight. This completes the proof of item (iii).
The proofs of items (i) and (ii) are similar, and so, we only give the proof of item (ii). First, suppose that ${\mathcal{G}}$ is tight. Let $\mathop{\rm conv}L_{q}(\mu,{\mathcal{H}}_{{\mathcal{G}}})$ be the convex hull of $L_{q}(\mu,{\mathcal{H}}_{{\mathcal{G}}})$, which is a subset of $L_{q}\bigl{(}\mu,{\mathcal{H}}_{{\mathcal{P}}}\bigr{)}$ as $L_{q}\bigl{(}\mu,{\mathcal{H}}_{{\mathcal{P}}}\bigr{)}$ is convex. Since $q\in(1,\infty)$, $L_{q}\bigl{(}\mu,{\mathcal{H}}_{k}\bigr{)}$ is reflexive. Therefore, weak-topology (i.e., kernel mean embedding topology) coincides with the weak-topology. Moreover, for convex sets, closedness with respect to the weak-topology and norm-topology are equivalent. Therefore, it is sufficient to prove that $\mathop{\rm conv}L_{q}(\mu,{\mathcal{H}}_{{\mathcal{G}}})$ is closed in norm-topology.

### v2

@@ 3.2 Two weak topologies revisited
3.2.1 $w^{*}$-topology (or Borkar topology) on Stochastic Kernels
In this section, we introduce the $w^{*}$-topology on the set of stochastic kernels, a concept developed in [20][19, Section 2.4] for randomized Markov policies. Later, in [63], $w^{*}$-topology is used to establish the existence of optimal policies in team decision problems. This topology has proven to be a valuable tool in analyzing such problems. We will begin by recalling its formal definition and then proceed to discuss some of its fundamental properties.
@@ 3.2.1 $w^{*}$-topology (or Borkar topology) on Stochastic Kernels
which turns $(C_{0}({\mathsf{U}}),\|\cdot\|)$ into a separable Banach space. Let $\|\cdot\|_{TV}$ denote the total variation norm on ${\mathcal{M}}({\mathsf{U}})$. The following result is a well-known Riesz-Markov-Kakutani representation theorem, which states that the topological dual of $C_{0}({\mathsf{U}})$ is ${\mathcal{M}}({\mathsf{U}})$.
@@
[33, Theorem 7.17] For any $\nu\in{\mathcal{M}}({\mathsf{U}})$ and $g\in C_{0}({\mathsf{U}})$, let $I_{\nu}(g)\triangleq\langle\nu,g\rangle$, where
@@
Furthermore, the norm on ${\mathcal{M}}({\mathsf{U}})$ induced by this duality is the total variation norm.
@@
$\|\gamma\|_{q}$ $=\begin{cases}\left(\int_{{\mathsf{Y}}}\|\gamma(y)\|^{q}_{TV}\,\mu(dy)\right)^{1/q},&\,\,\text{if}\,\,q<\infty\\ {\mathrm{ess}}\sup_{y\in{\mathsf{Y}}}\|\gamma(y)\|_{TV},&\,\,\text{if}\,\,q=\infty.\end{cases}$ (7)
Since $C_{0}({\mathsf{U}})$ is separable, $L_{p}\bigl(\mu,C_{0}({\mathsf{U}})\bigr)$ is also separable for $1\leq p<\infty$. By Theorem 2.1, the topological dual of $L_{p}\bigl(\mu,C_{0}({\mathsf{U}})\bigr)$ is $\mathcal{L}_{q}\bigl(\mu,{\mathcal{M}}({\mathsf{U}})\bigr)$; that is
@@
[38, Definition C.1] A mapping $\gamma:{\mathsf{Y}}\rightarrow{\mathcal{M}}({\mathsf{U}})$ is called a stochastic kernel from $\mathsf{Y}$} to $\mathsf{U}$} if, for all $D\in{\mathcal{U}}$, the mapping ${\mathsf{Y}}\ni y\mapsto\gamma(y)(D)\in\mathbb{R}$ is ${\mathcal{Y}}/{\mathcal{B}}(\mathbb{R})$-measurable and $\gamma(y)\in{\mathcal{P}}({\mathsf{U}})$ for all $y\in{\mathsf{Y}}$. Let $\Gamma$ denote the set of all stochastic kernels from $\mathsf{Y}$} to $\mathsf{U}$}.
@@
Next, consider any closed set $F\subset{\mathsf{U}}$. The indicator function $1_{F}$ can be approximated pointwise by the continuous and bounded functions $h_{n}(u)=\max(1-nd_{{\mathsf{U}}}(u,F),0)$, where $d_{{\mathsf{U}}}$ denotes the metric on $\mathsf{U}$} and $d_{{\mathsf{U}}}(u,F)=\inf_{y\in F}d_{{\mathsf{U}}}(u,y)$. Consequently, this implies that the mapping $x\mapsto\gamma(x)(F)\in\mathbb{R}$ is also $\mathcal{X}/\mathcal{B}(\mathbb{R})$-measurable for any closed set $F$ in $\mathsf{U}$}. Thus, we can conclude that $\gamma$ is a stochastic kernel by [16, Proposition 7.25].
@@
Let ${\mathcal{P}}_{{}_{\leq 1}}({\mathsf{U}})$ denote the set of sub-probability measures in ${\mathcal{M}}({\mathsf{U}})$. Then, since $\mathcal{L}_{q}\bigl(\mu,{\mathcal{P}}_{{}_{\leq 1}}({\mathsf{U}})\bigr)$ is proved to be closed in $\mathcal{L}_{q}\bigl(\mu,{\mathcal{M}}({\mathsf{U}})\bigr)$ with respect to $w^{*}$-topology (see [63]) and is a bounded subset, it is (sequentially) compact and metrizable with respect to $w^{*}$-topology by Banach-Alaoglu Theorem. Note that $\mathcal{L}_{q}\bigl(\mu,{\mathcal{P}}({\mathsf{U}})\bigr)$ is a subset of $\mathcal{L}_{q}\bigl(\mu,{\mathcal{P}}_{{}_{\leq 1}}({\mathsf{U}})\bigr)$, and so, we can endow $\mathcal{L}_{q}\bigl(\mu,{\mathcal{P}}({\mathsf{U}})\bigr)$ with relative $w^{*}$-topology. Unfortunately, $\mathcal{L}_{q}\bigl(\mu,{\mathcal{P}}({\mathsf{U}})\bigr)$ is not closed in this topology as shown below [^5] In [63], we established this result by constructing a counterexample specifically for the case when $\mathsf{U}$$\mathbb{R}$, which is relatively easy. In Lemma 3.8 above, we extend this by providing a counterexample applicable to any locally compact space $\mathsf{U}$}. This counterexample draws inspiration from the elegant proof presented in [67, Lemma 12]..
@@
First note that if $\mathsf{U}$} is compact, then $C_{0}({\mathsf{U}})=C_{b}({\mathsf{U}})=C({\mathsf{U}})$. Moreover, if $\gamma_{\lambda}\rightharpoonup^{*}\gamma$ with respect to $w^{*}$-topology, then $\mu\otimes\gamma_{\lambda}({\mathsf{Y}}\times\cdot)\rightarrow\mu\otimes\gamma({\mathsf{Y}}\times\cdot)$ in vague topology, and so, in weak convergence topology as $C_{0}({\mathsf{U}})=C_{b}({\mathsf{U}})=C({\mathsf{U}})$. Since $\mu\otimes\gamma_{\lambda}({\mathsf{Y}}\times\cdot)\in{\mathcal{P}}({\mathsf{U}})$ for all $\lambda$ and ${\mathcal{P}}({\mathsf{U}})$ is a closed subset of ${\mathcal{M}}({\mathsf{U}})$ under weak convergence topology, we have $\mu\otimes\gamma({\mathsf{Y}}\times\cdot)\in{\mathcal{P}}({\mathsf{U}})$. Hence, $\gamma\in\mathcal{L}_{q}\bigl(\mu,{\mathcal{P}}({\mathsf{U}})\bigr)$.
@@
Since $\varepsilon$ is arbitrary, we have $\int_{{\mathsf{U}}}g(u)\,\nu_{n}(du)\rightarrow 0$ as $n\rightarrow\infty$ for any $g\in C_{0}({\mathsf{U}})$. Hence, $\nu_{n}$ converges to 0, where $0(\,\cdot\,)$ denotes the degenerate measure on $\mathsf{U}$}; that is, $0(D)=0$ for all $D\in{\mathcal{U}}$.
@@
$=\int_{{\mathsf{Y}}}\lim_{n\rightarrow\infty}\int_{{\mathsf{U}}}f(y)(u)\,\nu_{n}(du)\,\mu(dy)\text{ (as $\|f(y)\|$ is $\mu$-integrable)}$
$=0\text{ (as $f(y)\in C_{0}({\mathsf{U}})$)}.$
@@ 3.2.2 The Young Narrow Topology on Stochastic Kernels
Conversely, let $h\in\mathop{\rm Car}_{0}({\mathsf{X}}\times{\mathsf{U}})$ and define $f_{h}:{\mathsf{X}}\rightarrow C_{0}({\mathsf{U}})$ as $f_{h}(x)=h(x,\,\cdot\,)$. Then, for any $\gamma\in{\mathcal{M}}({\mathsf{U}})$, the function ${\mathsf{X}}\ni x\mapsto\int_{{\mathsf{U}}}f_{h}(x)(u)\,\gamma(du)\in\mathbb{R}$ is measurable. Hence, $f_{h}$ is weakly measurable. But since $C_{0}({\mathsf{U}})$ is separable, $f_{h}$ is also strongly measurable by [56, Theorem 4.2.4-(c)]. ∎
@@
Let us now define Young narrow topology on $\mathcal{L}_{q}\bigl(\mu,{\mathcal{M}}({\mathsf{U}})\bigr)$ [56, Definition 4.7.11] [^6] Normally, Young narrow topology is defined only on the set of stochastic kernels $\mathcal{L}_{q}\bigl(\mu,{\mathcal{P}}({\mathsf{U}})\bigr)$. But, here, we extend this definition to the set of all weakly-measurable functions from $\mathsf{Y}$} to ${\mathcal{M}}({\mathsf{U}})$. Moreover, in the original formulation, the topology is induced by the set of $C_{b}$-Caratheodory functions that are in $L_{1}(\mu,C_{b}({\mathsf{U}}))$. Here, we extend this definition to $L_{p}(\mu,C_{b}({\mathsf{U}}))$ for any $1\leq p<\infty$.. Young narrow topology on $\mathcal{L}_{q}\bigl(\mu,{\mathcal{M}}({\mathsf{U}})\bigr)$ is the smallest topology for which the mapping
@@
In Young narrow topology, one can prove that $\mathcal{L}_{q}\bigl(\mu,{\mathcal{P}}({\mathsf{U}})\bigr)$ is a closed (as opposed to $w^{*}$-topology and weak kernel mean embedding topology) and metrizable subset of $\mathcal{L}_{q}\bigl(\mu,{\mathcal{M}}({\mathsf{U}})\bigr)$ [56, Proposition 4.7.14]. However, $\mathcal{L}_{q}\bigl(\mu,{\mathcal{P}}({\mathsf{U}})\bigr)$ is not relatively (sequentially) compact in this topology unless $\mathsf{U}$} is compact because of the lack of the duality result (i.e., The Banach-Alaoglu theorem is not applicable in this context).
@@ 3.3 Comparison of Weak Topologies
Interestingly, to achieve closure in the $w^{*}$-topology and the weak kernel mean embedding topology, a tightness condition is required. This condition effectively ensures uniform concentration of probability measures on compact sets, a necessity highlighted by counterexample in Lemma 3.8. Similarly, for the Young narrow topology to attain relative (sequential) compactness, the same tightness condition must be satisfied, as implied by Prokhorov’s theorem. Consequently, in terms of ensuring the existence of optimal solutions, all three topologies provide comparable benefits.
@@
The proofs of these results are almost identical, therefore, we only prove (i). To prove the forward direction of (i), let $\gamma_{\lambda}\rightharpoonup^{*}\gamma$ with respect to weak kernel mean embedding topology. Fix any $E\in{\mathcal{Y}}$. For any $h\in{\mathcal{H}}_{k}$, define the following simple function $f\triangleq h\,1_{E}$ in $L_{p}(\mu,{\mathcal{H}}_{k})$. Then, we have
@@
Hence, $\mu\otimes\gamma_{\lambda}(E\times\cdot)\rightarrow\mu\otimes\gamma(E\times\cdot)$ in weak MMD topology. Since $\mu\otimes\gamma_{\lambda}(E\times{\mathsf{U}})=\mu\otimes\gamma(E\times{\mathsf{U}})=\mu(E)$, by [67, Lemma 5 and Remark 6], $\mu\otimes\gamma_{\lambda}(E\times\cdot)\rightarrow\mu\otimes\gamma(E\times\cdot)$ in MMD topology. This completes the proof of the forward part.
For the converse part, for all $E\in{\mathcal{Y}}$, let $\mu\otimes\gamma_{\lambda}(E\times\cdot)\rightarrow\mu\otimes\gamma(E\times\cdot)$ in MMD topology. Consider any simple function $f\triangleq\sum_{i=1}^{k}h_{i}\,1_{E_{i}}$ in $L_{p}(\mu,{\mathcal{H}}_{k})$. Then, we have
@@
Note that for a fixed $E\in\mathcal{Y}$, the total masses of the sub-probability measures $\{\mu\otimes\gamma_{\lambda}(E\times\cdot)\}$ and $\mu\otimes\gamma(E\times\cdot)$ are both equal to $\mu(E)$ (assuming, without loss of generality, that $\mu(E)>0$). As used several times so far, by [67, Lemma 5], the MMD topology, vague topology, and weak convergence topology are equivalent for the set of probability measures. Since the aforementioned sub-probability measures have the same mass as $\mu(E)$, we can normalize them using the constant $\mu(E)$ to convert them into probability measures and apply the equivalence result from [67, Lemma 5]. This completes the proof. ∎
@@
In other words, tightness ensures that distributions in $\mathcal{G}$} uniformly concentrates on compact sets.
@@
where $\mu\otimes\gamma(dy,du)\triangleq\mu(dy)\,\gamma(y)(du)$. One can prove that ${\mathcal{C}}_{{\mathcal{G}}}$ is tight in the sense of the Definition 4.7.20 in [56]. Hence, by [56, Theorem 4.7.22], $\mathcal{L}_{q}(\mu,{\mathcal{G}})$ is relatively (sequentially) compact with respect to the Young narrow topology. Conversely, suppose that $\mathcal{L}_{q}(\mu,{\mathcal{G}})$ is relatively (sequentially) compact with respect to the Young narrow topology and $\mathsf{U}$} is complete. This implies that $\mathcal{G}$} is also relatively (sequentially) compact with respect to the weak topology. Indeed, let $\{\nu_{n}\}$ be a sequence in $\mathcal{G}$} and for each $n$, define the constant function $\gamma_{n}(y)=\nu_{n}$ for all $y\in{\mathsf{Y}}$. Since $\mathcal{L}_{q}(\mu,{\mathcal{G}})$ is relatively (sequentially) compact with respect to the Young narrow topology, there exists a subsequence $\{\gamma_{n_{k}}\}$ of $\{\gamma_{n}\}$ that converges with respect to the Young narrow topology. This implies that subsequence $\{\nu_{n_{k}}\}$ of $\{\nu_{n}\}$ also converges with respect to the weak topology. Hence, $\mathcal{G}$} is relatively (sequentially) compact with respect to the weak topology. By Prohorov’s theorem [39, Theorem 1.4.12], $\mathcal{G}$} is tight. This completes the proof of item (iii).
The proofs of items (i) and (ii) are somewhat similar. Since the proof of item (i) is also rather standard, we present only the proof of item (ii). First, suppose that $\mathcal{G}$} is tight. Let $\mathop{\rm conv}L_{q}(\mu,{\mathcal{H}}_{{\mathcal{G}}})$ be the convex hull of $L_{q}(\mu,{\mathcal{H}}_{{\mathcal{G}}})$, which is a subset of $L_{q}\bigl(\mu,{\mathcal{H}}_{{\mathcal{P}}}\bigr)$ as $L_{q}\bigl(\mu,{\mathcal{H}}_{{\mathcal{P}}}\bigr)$ is convex. Since $q\in(1,\infty)$, $L_{q}\bigl(\mu,{\mathcal{H}}_{k}\bigr)$ is reflexive. Therefore, weak-topology (i.e., kernel mean embedding topology) coincides with the weak-topology. Moreover, for convex sets, closedness with respect to the weak-topology and norm-topology are equivalent. Therefore, it is sufficient to prove that $\mathop{\rm conv}L_{q}(\mu,{\mathcal{H}}_{{\mathcal{G}}})$ is closed in norm-topology.